# GST-UNet: Spatiotemporal Causal Inference with Time-Varying Confounders

### v1

@@ Appendix A Extended Literature Review
We aim to show that under 3.1 and 3.2, the CAPOs in Equation 1 can be identified recursively from a single time series via a sequence of conditional expectations.
@@
We first demonstrate the recursive decomposition of the intractable expectation in the CAPO definition (Equation 1). While this expectation is theoretically well-defined, it cannot be directly estimated in practice due to the limited availability of data. Specifically, we only observe a single time series, meaning we have just one sample of the history at time $t+\tau$ for each $t$. Nevertheless, as we will show, we can convert these expectations into expectations over prefix-based segments that allow us to estimate these quantities from the data.
@@
$=\mathbb{E}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t},\mathbf{A}_{t}=\mathbf{a}_{t}]$ (Sequential ignorability and positivity (3.1))
@@
We now show how to estimate the nested expectations using the prefix data. First, by 3.2, we can rewrite the inner-most expectation as
@@
Thus, by using 3.1, we can write this expectation over the prefix data which we have many samples of. Now consider the next nested expectation:
@@
$=\int_{P}Q_{\tau}(\mathbf{h}_{1:t+\tau-1}^{\mathbf{a}},\mathbf{a}_{t+\tau-1})p(x_{t+\tau-1},y_{t+\tau-1}\mid\phi(\mathbf{h}^{\mathbf{a}}_{t+\tau-2},\mathbf{a}_{t+\tau-2}))d(x_{t+\tau-1},y_{t+\tau-1})$ (3.2)
@@
Appendix C Experimental Details
@@
In this appendix, we provide further information on the simulation experiments (Section 5.1) and the real-world wildfire application (Section 5.2), including exact parameter settings, model architecture and execution details, hyperparameter selection strategies, and validation procedures. All code for generating, preprocessing, and analyzing both the synthetic and real-world datasets—and for training and evaluating GST-UNet—is available at https://github.com/moprescu/GSTUNet, with step-by-step replication instructions in the repository’s README.md.
@@
For both applications, GST-UNet employs a U-Net backbone with a single ConvLSTM layer (hidden dimension 32) and a contracting-expanding path of channel sizes $16\rightarrow 32\rightarrow 64\rightarrow 128\rightarrow 256$. The G-computation heads are implemented as shallow feed-forward neural networks that operate on the U-Net feature maps at each grid cell for G-computation. In practice, to ensure stable ConvLSTM training and reduce computational overhead, we truncate the input history to a fixed length. All neural networks are implemented via the nn module in PyTorch [Paszke et al., 2019]. Experiments were conducted on an NVIDIA A100 (Ampere) GPU using the Perlmutter system at the National Energy Research Scientific Computing Center (NERSC). The synthetic experiments required roughly 55 minutes per hyperparameter set, while the wildfire experiment completed in about 5 minutes.
@@ C.1 Synthetic Experiments
$\mathbf{Y}_{t}$ $=\gamma_{0}+\gamma_{1}\bigl{(}{\color[rgb]{0.890625,0.1015625,0.109375}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.1015625,0.109375}K_{YA}*\mathbf{A}_{t-1}}\bigr{)}+\gamma_{2}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\tfrac{1}{L}\!\sum_{l=1}^{L}}\,\bigl{(}{\color[rgb]{0.0546875,0.40234375,0.890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0546875,0.40234375,0.890625}K_{YX}*\mathbf{X}_{t-l}}\bigr{)}$
$\qquad+\gamma_{3}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\mathbf{Y}_{t-1}}+\epsilon_{Y},$
@@
    $\gamma_{0}=2.0,\;\gamma_{1}=1.5,\;\gamma_{2}=0.5,\;\gamma_{3}=0.5,\;K_{YX}=\frac{1}{16}\begin{pmatrix}1.0&1.0&1.0\\ 1.0&8.0&1.0\\ 1.0&1.0&1.0\end{pmatrix},\;K_{YA}=\frac{1}{16}\begin{pmatrix}1.0&1.0&1.0\\ 1.0&8.0&1.0\\ 1.0&1.0&1.0\end{pmatrix}.$
@@
We use $L=5$ temporal lags for $\mathbf{X}$ and $\mathbf{Y}$, a seed of $42$ for reproducibility. See Figure 4 for representative $t=100$ snapshots of $X_{100}$, $A_{100}$, and $Y_{101}$ under varying $\beta_{1}$.
@@
Neural Architectures. The GST-UNet comprises a single ConvLSTM layer (hidden dimension $32$), followed by a U-Net with channel sizes $16\!\rightarrow\!32\!\rightarrow\!64\!\rightarrow\!128\!\rightarrow\!256$. Its G-computation heads are shallow feed-forward networks operating on the final U-Net feature maps at each grid cell; both the U-Net’s output dimension ($d_{h}$) and the G-head architecture (number of layers, hidden size) are treated as hyperparameters. The UNet+ baseline uses the same ConvLSTM+U-Net backbone as GST-UNet but outputs a single channel ($d_{h}=1$), omitting any G-computation. For direct comparison, we also implement STCINet [Ali et al., 2024] with an identical ConvLSTM+U-Net backbone, and retaining their original Latent Factor Model (LFM) details.
@@
Training Details. We randomly initialize all model parameters (GST-UNet and baselines) with Xavier uniform weights [Glorot and Bengio, 2010]. We use the Adam optimizer [Kingma, 2014] with an initial learning rate, halving it whenever the validation loss plateaus for a specified scheduler patience. To mitigate overfitting, we adopt early stopping when the validation loss fails to improve for a specified early stopping patience epochs. Validation uses 40 of the 190 training prefixes, and the total training is capped at 100 epochs. We tune the following hyperparameters: (i) batch size, learning rate, scheduler patience, and early stopping patience (common to all models); (ii) for GST-UNet, the curriculum period and learning rate for curriculum phases, the U-Net output dimension $d_{h}$, and the number and width of hidden layers in the feed-forward G-heads. Table 2 lists the hyperparameter ranges considered, with the values yielding the best validation performance in bold.
[…]
C.2 Wildfire Application

### v2

@@ Appendix B Proof of Theorem 1
We aim to show that under Assumption˜1 and Assumption˜2, the CAPOs in Equation˜1 can be identified recursively from a single time series via a sequence of conditional expectations.
@@
We first demonstrate the recursive decomposition of the intractable expectation in the CAPO definition (Equation˜1). While this expectation is theoretically well-defined, it cannot be directly estimated in practice due to the limited availability of data. Specifically, we only observe a single time series, meaning we have just one sample of the history at time $t+\tau$ for each $t$. Nevertheless, as we will show, we can convert these expectations into expectations over prefix-based segments that allow us to estimate these quantities from the data.
@@
$=\mathbb{E}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t},\mathbf{A}_{t}=\mathbf{a}_{t}]$ (Sequential ignorability and positivity (Assumption 1))
@@
We now show how to estimate the nested expectations using the prefix data. First, by Assumption˜2, we can rewrite the inner-most expectation as
@@
Thus, by using Assumption˜1, we can write this expectation over the prefix data which we have many samples of. Now consider the next nested expectation:
@@
$=\int_{P}Q_{\tau}(\mathbf{h}_{1:t+\tau-1}^{\mathbf{a}},\mathbf{a}_{t+\tau-1})p(x_{t+\tau-1},y_{t+\tau-1}\mid\phi(\mathbf{h}^{\mathbf{a}}_{t+\tau-2},\mathbf{a}_{t+\tau-2}))d(x_{t+\tau-1},y_{t+\tau-1})$ (Assumption 2)
@@
Appendix C Consistency of the Iterative G-Computation Estimator
In this section, we state the conditions under which the iterative G-computation procedure in Section˜4.2 yields a consistent estimator, and show that our implementation of the $Q_{k}$ estimators satisfies these conditions.
Notation: We denote the $L_{2}$ norm of a function $f$ as $\|f\|_{2}:=\mathbb{E}_{P}[f(X)^{2}]^{1/2}$, where the expectation is over the probability distribution $P$. The notation $\widehat{f}_{n}$ represents the estimated value of a parameter or function learned on $n$ data points, where $f$ is the true value. For a sequence of random variables $\{Z_{n}\}_{n\geq 1}$ we write $Z_{n}=o_{p}(1)$ if $\Pr(|Z_{n}|>\varepsilon)\to 0$ for every $\varepsilon>0$, i.e. $Z_{n}\xrightarrow{p}0$.
To begin, we introduce the following stochastic equicontinuity condition from [43]:
Definition 1 (Stochastic equicontinuity [43, Def. 1.5.7]).
Let $(\mathcal{Z},d)$ be a semi-metric space and $\{\widehat{f}_{n}\}_{n\geq 1}\subset\ell^{\infty}(\mathcal{Z})$ a sequence of random functions. It is uniformly stochastically equi-continuous if, for every $\epsilon>0,\eta>0$, there exists a $\delta>0$ such that
$\limsup_{n\to\infty}\;P\!\Bigl(\sup_{d(z,z^{\prime})\leq\delta}\bigl|\widehat{f}_{n}(z)-\widehat{f}_{n}(z^{\prime})\bigr|>\epsilon\Bigr)<\eta.$
Stochastic equicontinuity ensures that, with high probability, each estimator changes only slightly when its input is perturbed by a small amount. It is strictly weaker than global Lipschitz continuity – any family that is Lipschitz on a bounded domain with constants bounded in probability automatically satisfies Definition 1. We impose this condition in Theorem 3 so that the $o_{p}(1)$ error in the learned embedding propagates to only $o_{p}(1)$ errors in the G-heads, making the recursive estimator consistent.
The following theorem restates Theorem 2 from the main text in full detail and provides its proof.
Theorem 3 (Consistency under Uniform Stochastic Equicontinuity).
Suppose the conditions of Theorem 1 hold, and let $\widehat{\phi}$ be a learned embedding. Define $\textbf{Z}_{k}:=(\textbf{H}_{1:t+k},\textbf{A}_{t+k})$, and recursively define the learned estimators $\widehat{Q}_{k}(\textbf{Z}_{k-1};\widehat{\phi}):=\widehat{\mathbb{E}}_{\mathbf{P}}[\widehat{Q}_{k+1}(\textbf{Z}_{k};\widehat{\phi})\mid\widehat{\phi}(\textbf{Z}_{k})]$ for $k=1,\dots,\tau$, with terminal condition $\widehat{Q}_{\tau+1}(\textbf{Z}_{\tau};\widehat{\phi})=Y^{t+\tau}$. Assume that $\{\widehat{Q}_{k}\}_{k=1}^{\tau}$ are obtained via the iterative G-computation algorithm. If:
  1. (i)
    $\|\widehat{\phi}-\phi\|_{2}=o_{p}(1)$;
  2. (ii)
    $\|\widehat{Q}_{k}\!\bigl(\textbf{Z}_{k-1};\phi\bigr)-Q_{k}\!\bigl(\textbf{Z}_{k-1};\phi)\bigr)\|_{2}=o_{p}(1)$ for all $k$;
  3. (iii)
    for every $k$ the random maps $z\mapsto\widehat{Q}_{k}(h,a;z)$ are stochastically equicontinuous on $\operatorname{Im}\phi$ (Definition 1), and $Q_{k}(\cdot)$ is uniformly continuous there,
then
$\left\|\widehat{Q}_{1}\!\bigl(\textbf{Z}_{0};\widehat{\phi}\bigr)-Q_{1}\!\bigl(\textbf{Z}_{0};\phi\bigr)\right\|_{2}=o_{p}(1).$
Thus the recursive G-computation estimator is (probabilistically) consistent.
Proof.
We proceed by reverse induction on $k$, starting from $k=\tau$ and working backward to $k=1$. For each $k$, we aim to show:
$\|\widehat{Q}_{k}(\mathbf{Z}_{k-1};\widehat{\phi})-Q_{k}(\mathbf{Z}_{k-1};\phi)\|_{2}=o_{p}(1).$
Base case ($k=\tau$). By definition, $\widehat{Q}_{\tau+1}(\mathbf{Z}_{\tau};\widehat{\phi})=Y^{t+\tau}$, which is observed. Thus,
$\widehat{Q}_{\tau}(\mathbf{Z}_{\tau-1};\widehat{\phi})=\widehat{\mathbb{E}}_{\mathbf{P}}[Y^{t+\tau}\mid\widehat{\phi}(\mathbf{Z}_{\tau})]\quad\text{and}\quad Q_{\tau}(\mathbf{Z}_{\tau-1};\phi)=\mathbb{E}[Y^{t+\tau}\mid\phi(\mathbf{Z}_{\tau})].$
We decompose the difference:
$\left\|\widehat{Q}_{\tau}(\mathbf{Z}_{\tau-1};\widehat{\phi})-Q_{\tau}(\mathbf{Z}_{\tau-1};\phi)\right\|_{2}\leq\underbrace{\left\|\widehat{Q}_{\tau}(\mathbf{Z}_{\tau-1};\widehat{\phi})-\widehat{Q}_{\tau}(\mathbf{Z}_{\tau-1};\phi)\right\|_{2}}_{\Lambda_{1}}+\underbrace{\left\|\widehat{Q}_{\tau}(\mathbf{Z}_{\tau-1};\phi)-Q_{\tau}(\mathbf{Z}_{\tau-1};\phi)\right\|_{2}}_{\Lambda_{2}}.$
Term $\Lambda_{2}$ is $o_{p}(1)$ by assumption (ii). Term $\Lambda_{1}$ converges to zero in probability due to (i) $\|\widehat{\phi}-\phi\|_{2}=o_{p}(1)$ and (iii) stochastic equicontinuity of $\widehat{Q}_{\tau}$. Therefore,
$\left\|\widehat{Q}_{\tau}(\mathbf{Z}_{\tau-1};\widehat{\phi})-Q_{\tau}(\mathbf{Z}_{\tau-1};\phi)\right\|_{2}=o_{p}(1).$
Inductive step. Suppose for some $k+1\leq\tau$ that
$\left\|\widehat{Q}_{k+1}(\mathbf{Z}_{k};\widehat{\phi})-Q_{k+1}(\mathbf{Z}_{k};\phi)\right\|_{2}=o_{p}(1).$
We now consider
$\widehat{Q}_{k}(\mathbf{Z}_{k-1};\widehat{\phi})=\widehat{\mathbb{E}}_{\mathbf{P}}[\widehat{Q}_{k+1}(\mathbf{Z}_{k};\widehat{\phi})\mid\widehat{\phi}(\mathbf{Z}_{k})],\quad Q_{k}(\mathbf{Z}_{k-1};\phi)=\mathbb{E}[Q_{k+1}(\mathbf{Z}_{k};\phi)\mid\phi(\mathbf{Z}_{k})].$
Again, decompose:
$\left\|\widehat{Q}_{k}(\mathbf{Z}_{k-1};\widehat{\phi})-Q_{k}(\mathbf{Z}_{k-1};\phi)\right\|_{2}$ $\leq\left\|\widehat{Q}_{k}(\mathbf{Z}_{k-1};\widehat{\phi})-\widehat{Q}_{k}(\mathbf{Z}_{k-1};\phi)\right\|_{2}$
$\quad+\left\|\widehat{Q}_{k}(\mathbf{Z}_{k-1};\phi)-Q_{k}(\mathbf{Z}_{k-1};\phi)\right\|_{2}.$
The second term is $o_{p}(1)$ by assumption (ii). The first term is also $o_{p}(1)$ because $\widehat{\phi}\to\phi$ in $L_{2}$ and the stochastic equicontinuity of $\widehat{Q}_{k}$ ensures that perturbations in $\phi$ yield small changes in predictions uniformly over $\operatorname{Im}\phi$. Thus,
$\|\widehat{Q}_{k}(\mathbf{Z}_{k-1};\widehat{\phi})-Q_{k}(\mathbf{Z}_{k-1};\phi)\|_{2}=o_{p}(1).$
By induction, the result holds for all $k=\tau,\tau-1,\dots,1$, and in particular:
$\|\widehat{Q}_{1}(\mathbf{Z}_{0};\widehat{\phi})-Q_{1}(\mathbf{Z}_{0};\phi)\|_{2}=o_{p}(1).$
Thus, the proof is now complete. ∎
Example 1 (Feed-forward or convolutional heads).
Suppose each G-computation head $Q_{k}(\cdot;z)$ is implemented as a depth-$d$ neural network
$\Psi(z)=W_{d}\sigma_{d-1}\bigl(\cdots\sigma_{1}(W_{1}z)\bigr),$
where the activations $\sigma_{\ell}$ are Lipschitz continuous (e.g., ReLU, Leaky ReLU, SoftPlus, Tanh, Sigmoid, or ArcTan). If each layer weight satisfies a spectral norm bound $\|W_{\ell}\|_{2}\leq\rho_{\ell}<\infty$, then $\Psi$ is globally Lipschitz on $\mathbb{R}^{h}$ with constant $L=\prod_{\ell}\rho_{\ell}$, and thus uniformly continuous on any compact subset. This implies the stochastic equicontinuity condition in Definition˜1.
In practice, norm control can be enforced via weight decay, spectral normalization, or weight clipping during training. Similarly, the encoder output $\widehat{\phi}(H,A)$ can be bounded—e.g., through normalization or clipping—so its image lies in a compact subset of $\mathbb{R}^{h}$. Together, these ensure the continuity and equicontinuity conditions required by Theorem˜3.
The same argument applies to convolutional networks, since 2-D convolutions are linear operators whose induced matrix representations also admit spectral norm bounds controlled via spectral normalization.
Appendix D Experimental Details
@@
In this appendix, we provide further information on the simulation experiments (Section˜6.1) and the real-world wildfire application (Section˜6.2), including exact parameter settings, model architecture and execution details, hyperparameter selection strategies, and validation procedures. All code for generating, preprocessing, and analyzing both the synthetic and real-world datasets—and for training and evaluating GST-UNet—is available at https://github.com/moprescu/GSTUNet, with step-by-step replication instructions in the repository’s README.md.
@@
For both applications, GST-UNet employs a U-Net backbone with a single ConvLSTM layer (hidden dimension 32) and a contracting-expanding path of channel sizes $16\rightarrow 32\rightarrow 64\rightarrow 128\rightarrow 256$. The G-computation heads are implemented as shallow feed-forward neural networks that operate on the U-Net feature maps at each grid cell for G-computation. In practice, to ensure stable ConvLSTM training and reduce computational overhead, we truncate the input history to a fixed length. All neural networks are implemented via the nn module in PyTorch [32]. Experiments were conducted on an NVIDIA A100 (Ampere) GPU using the Perlmutter system at the National Energy Research Scientific Computing Center (NERSC). The synthetic experiments required roughly 55 minutes per hyperparameter set, while the wildfire experiment completed in about 5 minutes.
@@ D.1 Synthetic Experiments
$\mathbf{Y}_{t}$ $=\gamma_{0}+\gamma_{1}\bigl({\color[rgb]{0.890625,0.1015625,0.109375}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.1015625,0.109375}K_{YA}*\mathbf{A}_{t-1}}\bigr)+\gamma_{2}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\tfrac{1}{L}\!\sum_{l=1}^{L}}\,\bigl({\color[rgb]{0.0546875,0.40234375,0.890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0546875,0.40234375,0.890625}K_{YX}*\mathbf{X}_{t-l}}\bigr)+\gamma_{3}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\mathbf{Y}_{t-1}}+\epsilon_{Y},$
@@
    $\gamma_{0}=2.0,\;\gamma_{1}=1.5,\;\gamma_{2}=0.5,\;\gamma_{3}=0.5$
    $K_{YX}=\frac{1}{16}\begin{pmatrix}1.0&1.0&1.0\\ 1.0&8.0&1.0\\ 1.0&1.0&1.0\end{pmatrix},\;K_{YA}=\frac{1}{16}\begin{pmatrix}1.0&1.0&1.0\\ 1.0&8.0&1.0\\ 1.0&1.0&1.0\end{pmatrix}$.
@@
Table 4: Ablation on spatial kernel size ($\tau=5$). Removing neighbor aggregation ($1\times 1$ kernel) degrades performance, confirming the need to model spatial spill-overs.
Kernel size $\beta_{1}{=}0.0$ $0.5$ $1.0$ $1.5$ $2.0$
$3\times 3$ 0.33 $\pm$ 0.004 0.35 $\pm$ 0.004 0.40 $\pm$ 0.005 0.44 $\pm$ 0.004 0.40 $\pm$ 0.005
$1\times 1$ 0.53 $\pm$ 0.004 0.55 $\pm$ 0.005 0.54 $\pm$ 0.005 0.60 $\pm$ 0.007 0.64 ± 0.006
We use $L=5$ temporal lags for $\mathbf{X}$ and $\mathbf{Y}$, a seed of $42$ for reproducibility. The parameter values were chosen such that the simulation remains stable (i.e., the process does not diverge). See Figure 4 for representative $t=100$ snapshots of $X_{100}$, $A_{100}$, and $Y_{101}$ under varying $\beta_{1}$.
@@
Neural Architectures. The GST-UNet comprises a single ConvLSTM layer (hidden dimension $32$), followed by a U-Net with channel sizes $16\!\rightarrow\!32\!\rightarrow\!64\!\rightarrow\!128\!\rightarrow\!256$. Its G-computation heads are shallow feed-forward networks operating on the final U-Net feature maps at each grid cell; both the U-Net’s output dimension ($d_{h}$) and the G-head architecture (number of layers, hidden size) are treated as hyperparameters. The UNet+ baseline uses the same ConvLSTM+U-Net backbone as GST-UNet but outputs a single channel ($d_{h}=1$), omitting any G-computation. For direct comparison, we also implement STCINet [1] with an identical ConvLSTM+U-Net backbone, and retaining their original Latent Factor Model (LFM) details.
IPWUNet Baseline. We adapt the Inverse Propensity Weighting (IPW) estimator from [49] to the spatiotemporal setting. Given estimated propensities $\hat{\pi}(\mathbf{a}_{l}\mid\mathbf{H}_{1:l})$, the estimator is defined as:
$\hat{Y}_{t+\tau}^{\text{IPW}}=\left(\prod_{l=t}^{t+\tau}\frac{\mathbb{I}[\mathbf{A}_{l}=\mathbf{a}_{l}]}{\hat{\pi}(\mathbf{a}_{l}\mid\mathbf{H}_{1:l})}\right),\quad\text{CAPO}=\mathbb{E}[\hat{Y}_{t+\tau}^{\text{IPW}}\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t}].$
We implement the IPWUNet baseline by reusing the UNet+ architecture (U-Net + ConvLSTM + Attention) for both propensity estimation and outcome prediction. Specifically, we first train the propensity model with a binary cross-entropy loss to estimate $\hat{\pi}(\mathbf{A}_{t}\mid\mathbf{H}_{t})$ at each time $t$. We then freeze this model and use the estimated weights to train a second instance of the same architecture with a weighted MSE loss, where pseudo-outcomes are reweighted by the estimated inverse propensities along the counterfactual treatment path. While this allows partial adjustment for time-varying confounding, the method does not correct for spatial interference and is sensitive to small propensity values, which can lead to high variance.
@@
Training Details. We randomly initialize all model parameters (GST-UNet and baselines) with Xavier uniform weights [17]. We use the Adam optimizer [21] with an initial learning rate, halving it whenever the validation loss plateaus for a specified scheduler patience. To mitigate overfitting, we adopt early stopping when the validation loss fails to improve for a specified early stopping patience epochs. Validation uses 40 of the 190 training prefixes, and the total training is capped at 100 epochs. We tune the following hyperparameters: (i) batch size, learning rate, scheduler patience, and early stopping patience (common to all models); (ii) for GST-UNet, the curriculum period and learning rate for curriculum phases, the U-Net output dimension $d_{h}$, and the number and width of hidden layers in the feed-forward G-heads. Table 3 lists the hyperparameter ranges considered, with the values yielding the best validation performance in bold.
Evaluation Procedure. We evaluate each model by averaging the root mean square error (RMSE) of the estimated CAPOs against ground truth across 50 test trajectories. Table 1 in the main text reports RMSE $\pm$ standard deviation for horizon lengths $\tau\in\{5,\,10\}$ and $\beta_{1}\in\{0,0.5,1.0,1.5,2.0\}$.
Table 5: Effect of increasing $T$ on RMSE for $\beta_{1}=2.0$. GST-UNet improves with more data, while baselines remain biased.
Model T=100 T=200 T=400 T=600 T=800
UNet+ 0.78 0.81 0.82 0.95 0.87
STCINet 0.80 0.90 1.04 1.02 0.91
GST-UNet 0.69 0.40 0.32 0.32 0.36
Effect of Varying $T$. We ran additional simulations varying the trajectory length $T\in\{100,200,400,600,800\}$ and $\beta_{1}=2.0$ while keeping the grid size fixed ($d_{x}=d_{y}=64$). Results are shown Table˜5. GST-UNet consistently improves with more data, while the baselines remain biased—even as $T$ increases. This highlights the importance of adjusting for time-varying confounding: without it, there is a persistent asymptotic bias.
Effect of Neighbor Aggregation. To evaluate the importance of spatial spill-over modeling, we ablate the convolutional kernel used in the ConvLSTM encoder. Table˜4 compares GST-UNet with a standard $3\times 3$ kernel against a variant that removes neighbor aggregation by using a $1\times 1$ kernel. Across all levels of confounding strength ($\beta_{1}$), performance deteriorates markedly when neighbor information is excluded, with RMSE increasing by 30–40%. This confirms that explicitly aggregating information from nearby locations is essential for capturing spatial interference and achieving unbiased counterfactual estimates.
D.2 Wildfire Application